OpenStudy (anonymous):

During a vacation, you go to Dinner on a rotating restaurant that is circular and has a radius of 40 ft. At 6:42pm, you take a seat at the window table. You finish dinner at 8:28 pm. What angle did your position rotate throughout your stay? How many feet did your position revolve?

OpenStudy (anonymous):

Please help! Or just give me an idea where to start?

OpenStudy (anonymous):

Need rotation rate info.

OpenStudy (anonymous):

It makes 1 rotation per hour.

OpenStudy (anonymous):

Fine. 1 rotation per hour is 360o per hour. 8:28-6:42= 106 minutes= 106/60 = 1.77 hr 1 rotation returns you where you were rotation moves you 077(360)=277at that, about 3/4 around

OpenStudy (anonymous):

Thank you!